Site hosted by Build your free website today!
Excerpts for Bravo Co. from Declassified S2/S3 Daily Journals of 1st Battalion, 20th Infantry, 11th Brigade 30 March, 04-14 May 1970

30 March 1970 (Webmaster's Note: 36 ele is 3rd Sqd/3rd Plt)
1115-B Co. The 36 ele engaged 16 NVA/VC dressed in green uniforms and carrying packs and 3 to 4 weapons grid 777388. They were 500 meters away when engaged. Gunships were requested. Shark 7 is on station at 1125. The 32 ele is standing by for C/A. RTO Roger. B32 and 12 ele's will C/A. The 12 ele from grid 783433 to 778395 and 32 ele from 789417 to 765390. Complete at 1225. RTO, roger.
1245. B Co. The 36 ele found one KIA, 25 to 30 years old and dressed in blue uniform with a ammo belt, at 778387. He was killed by small arms fire. He had a picture, M26 grenade and a US ammo pouch. The 12 ele found 12 to 15 bushels of potatoes and a small amount of rice. RTO,Rgr.

4 May 70
0808-B/39 ele provides road security at 813398 and the 169PF's sweep thru the area of contact and negative findings.
0900-B/35 is at 822398.
0943-B/38 ele is road security at 813398.
1500-Bde documents captured from a NVA officer on 03 May indicate attacks will begin on the night of 8 May in Quang Ngai Prov. The attacks will be made by IF, MF and NVA units. ARVN 4-4 in the Son Ve Valley captured the documents. The intel was passed by secure means to all units.
1630-B ele C/A to grid 790488.
1920-B Co. The 36 snk is at 807395.

05 May 70 (Death of Captain Lisowski)
0620-B Co. 36 ele. is moving to grid 780390 for final extraction security.
0630- The 57 ele is at cp1 at grid 775437 and the 56 ele is moving to 779452.
0643-57 ele is at cp2 at 774431 and 56 is at 779452.
0650-56 ele has closed with negative findings.
0710-56 and 57 elements have closed on 776446.
0717-We request an urgent dustoff for three US with multiple frag wounds from a booby trap at 788443. We also have one US KIA. The 74 ele, who was walking point, and hit a booby trapped hand frag with a trip wire attached and 74 received frag wounds to legs. The 31 ele was moving up to the 74 ele location and hit another booby trapped hand frag with a trip wire attached. The CO was carrying exploseives and blasting caps when the booby trap went off causing a secondary explosion. The CO was killed and 2 US WIA as a result. Line Numbers of the WIA are 29, 87, and 127. The KIA Line No. is 01. The CO was carrying 5 pounds of TNT, 4 to 6 blasting caps, and some time fuze and det cord. Dust off complete 0737.
(The 3 WIA were: Sgt. Darold D. Marcus; SP4 Larry J. Ogdan; PFC Richard S. Pina)
0800-B Co at 790445 found an 8 inch round and request EOD to blow it.
0845-B 74 and 31 ele closed and 36 ele is moving with 39 ele at cp2
0850-B 36 ele closed on 783433.
0935-79 and 39 eles have closed on 796395 and 784344 respectively.
1220-Div. graves registration at Bravo and 76 ele moving to 788443.
1415-76 ele closed on 783433 and Graves Registration had negative findings.
1745-Some VN females told 36 ele that 2 VC working in area 770380 posing as bamboo cuttters.
2040-The RF's are at 800397 and 36 ele is at 806395.

06 May 1970-BMNT@0530;Sunrise@0619;Sunset@1904;EENT@1954;Moonrise@0616;Moonset@1946 02% illum
0535-B Co's 37 ele has closed on 783433. RTO roger (Arty)
0638-B Co 58 ele is moving to 757457. RTO roger (Arty)
0642-B Co 76 ele is moving to 791448. RTO roger (Arty)
0705-B Co 57 ele closed on 783433.
0755-B Co The 39 ele is moving to 782417 and the 36,37, and 76 elements closed on 783433.RTO roger
0800-B Co The 58 ele has closed on 783433. RTO.roger(Arty)
0805-B Co The 38 ele is moving to 820395. RTO roger(Arty)
0855-B Co the 35 and 36 eles are on road security. RTO roger(Arty)
0940-B Co The 39 ele has closed on 783433 with neg findings. RTO roger
1340-B Co A VN child brought one M16E3 bouncing betty mine. It was on the red ball road at 778444 about 300 meters south of the access road. The mine was armed and ready to blow The mine is now at the Twin Villes and will be picked up by S2 and S5. RTO, roger.
1540-B Co A VN female reports that there is a part of a body in the vicinity of Van Truong. The 74 ele is moving out to check the area. The 74 ele requests a body bag for the left leg of a US (CPT Lisowski).RTO,roger
1653-B Co the 74 ele is moving to 795427 and the 34 ele is moving to 796407. RTO, roger.
1745-B Co the 74 ele has closed on 783433. RTO, roger (Arty) 1745-B Co The 56 ele is at 767450 where they engaged one VC with a weapon. It was an unknown type of weapon and the clothing was unidentifiable because they were 100 meters away. He evaded to west into a treeline. The 56 ele checked out the area with negative findings. RTO,roger
1748-B Co the 79 ele is at cp1 at 795427. RTO roger
1821 B Co the 79 ele is at cp2 at 791435. RTO roger
1835 B Co We request an urgent dustoff for one US with a frag wound to the shoulder resulting from a hand frag thrown by a VC from approx 10 feet. The 56 ele engaged the VC resulting in one VC KIA, one M2 Carbine CIA and 2 M26 grenades CIA. THe VC was dressed in black pants and green shirt. RTO, roger
NOTE FROM WEBMASTER: WIA was Sgt Melvin B. Welch.
1900-B Co The 79 ele spotted two VC approx 300 meters to their front (north) with a weapon. 81mm was called in on the area. It is unknown what type of weapon it was or the type of clothing which was worn. They evaded to the east. The area of 798442 was checked with negative findings. RTO, roger
1914-B Co We request a priority dustoff for one VC civilian who was shot in the back by a carbine. He was carried back to 771461 and could possibly be a VC. Completed at 1925. RTO, roger.
1920- B Co The 56 and 79 elements have closed. RTO, roger
1920-1/20th We request 2 platoons of E Troop,1-1 Cav, be OPCON to 1-20 for 5 days beginning 071300May70. The purpose is to search and clear. We also requested 2 flame tracks from 1-1 Cav from Chu Lai to work the Gaza Strip for 5 days. CPT Robinson turned down the request.
1930-B Co The 78 ele is moving to 789433.RTO,roger
1940-B Co The 56 ele VC KIA also had a NVA flag, a US poncho liner, an AK47 magazine(empty), a US pistol belt, two electrical blasting caps, 2 books (one was a plan to make a booby trap),one homemade booby trap (frag type), an ID with the name Pham Van Dam (VC type) issued by Trong Tan, and he had one letter addressed to a person in Saigon asking for money (Name:Trang). RTO, roger.
2020-B Co the 78 ele is at 789433, the 36 ele is at 799398, the 58 ele is at 772455 and the 39 ele is at 783422. RTO, roger(Arty)
2250-1-20th The Bn Commander talked to Bde S3 reference the Secret Message of increased activity. Bde S3 stated that no implementing instructions would be issued reference the Secret Message. Bde S3 stated that 1-20 Inf operations would be carried out as planned on 7 May 1970 and were satisfactory. No change in planned operations was necessary. The purpose of the call by the Bn Commander was to discuss redisposition of 1-20th Inf elements in response to subject message. 2330-1-20th The Secret message was passed on to A CO, B CO and C Co by CPT Bassett.

07 May 70
0745-B/56,78,36,39 eles have closed.
0805-B/39 to 774424 and 38 ele to 825398 for road security.
0821-B/57 closed on 783433.
1020-B/39 closed on 796396.
1500-1/20 All units informed of Rules of Engagement.
1818-RF ambush grid 798374 2 VC KIA.
2040-The PF snks are at 805386 and 766404.
2134-B Co informed of 10-15 stationary personnel sighted by radar at 781463.(br>

08 May 70
0013-B/54 1 squad VC and 3 sappers penetrated perimeter 776446 with shaped charges and grenades at Twin Villes, 54 ele took small arms and B40 rocket fire.
0050-B/54 ele received small arms fire and B40 rockets from Kim Giao. B CP received one B40 rocket.
0215-B CP received small arms fire from Van Truong 1. 1 US KIA, 3 US WIA, line nos. 8, 36, and 39. 2 VC KIA and 5 PFs WIA.
0400-VC broke off attack.
0455-TMF sighted unknown no. of personnel with metal detected at 795371.
0455-B Co. We request urgent dustoff for 5 PF with multiple frag wounds. 3 US WIA with minor frag wounds. ETA 10-15 min. complete at 0520.
0630-B Co. requests urgent dustoff for 2 Viet civilians with gunshot and frag wounds as result of B54 contact.
0843-B/38 ele road security at 825398.
1710 thru 1904-B/76,79,38,57 elements at grid 783433 with neg findings.
1940-B/36 snk is at 798398.
2000-B/36 ele is at 798398.
2045-B/56 ele canceled because the PFs would not go out because of a ball game tomorrow.

09 May 70
0743-B/37,39,36 elements closed grid 783433.
0750-B/38 ele moving to grid 825398 for road security.
0755-1/20 Bn. Tonight have out max ambushes from 718428 to753465;720420-778406 and swing south in arc.
1100-1/20. All units review Rules of Engagement.
1101-B/34 and 35 closed grid 783433.
1147-B/58 closed 783433.
1805-B Co perimeter at Twin Villes, block C4 from fougasse missing, trip-flare cut, perimeter wire cut, claymore missing.
1953-B/76 ele moving to NDP at 783436.
2030-B/76 closed due to sick personnel.
2350-B/35 ele at 169 RF Compound received 5 B40 rockets and approx 80-90 rounds of automatic weapons fire to their southwest at approx 150-200 meters. Suspected enemy grid is 785424 with neg casualties. Arty and 4.2 mortars fired.

10 May 70
0023-B/35 at 169 RF Compound dustoff 4 ARVN casualties with frag wounds, one baby with head wound.
0543-B/76 closed grid 7873433. 0840-B Co at grid 789417 two US WIA B40 frag wounds line nos. 61 and 67.
1025-B/95 ele found a dud 105 arty round at 782431 with Comp B end carved out and two M79 rounds stuffed in hole.
1210-B/38 ele moving to 825398.
1610-B/38 ele on road security and has now closed.
2158-B/36 ele at Duc Pho Compound took small arms fire from SE 600-800 meters away. The 169 RFs spotted 7 to 14 VC. Engaged with small arms, arty. Neg results.
2240-1/20 A new NVA L-17 Bn at grid 721455 with mission to attack LZ Liz (755432) and FSB Bronco (818385) from now to 22 May 70. They have 2-105 arty and resupplied today.
2320-B CP took 4 incoming rounds from 775430 and 790433. Fired 81mm.

11 May 70
0737-B/38 moving to 822398.
1050-B/74 engaged with 5 VC vic. 787438, urgent dust off for 1 US WIA head.
2005-B/79 cancelled due to lack of PF/RF

12 May 70 0921-B/36 ele at road security at 822394
1850-B/36 closed grid 783433.
2050-B Co. The ARVNs have snk locations at 795398 and 805397.

13 May 70 1952-B/38 snk is at 7923987.
2130-B/79 ele patrol for tomorrow is cancelled due to ops.
2255-B/38 closed on CP.

14 May 70 0647-B/38 and 57 elements moving out to 796389 and 773445, respectively.
0706-B/38 ele at cp1 at 796389.
0730-B/38 at cp2 at 801497.
1700-11 Bde. Black Aces Arty plane spotted 50 new fighting pos 3 NVA grid 7851.
2026-B/36 Tomorrow with RF will move to 795372 at 0400.
2330-The 58 ele patrol is cancelled.
Top of Page

Click for Main Page